# A Generalized Difference-in-Differences Estimator for Unbiased Estimation of Desired Estimands from Staggered Adoption and Stepped-Wedge Settings

### v1

@@ Abstract
Staggered treatment adoption arises in the evaluation of policy impact and implementation in a variety of settings. This occurs in both randomized stepped-wedge trials and non-randomized quasi-experimental designs using causal inference methods based on difference-in-differences analysis. In both settings, it is crucial to carefully consider the target estimand and possible treatment effect heterogeneities in order to estimate the effect without bias and in an interpretable fashion. This paper proposes a novel non-parametric approach to this estimation for either setting. By constructing an estimator using two-by-two difference-in-difference comparisons as building blocks with arbitrary weights, the investigator can select weights to target the desired estimand in an unbiased manner under assumed treatment effect homogeneity, and minimize the variance under an assumed working covariance structure. This provides desirable bias properties with a relatively small sacrifice in variance and power by using the comparisons efficiently. The method is demonstrated on toy examples to show the process, as well as in the re-analysis of a stepped wedge trial on the impact of novel tuberculosis diagnostic tools. A full algorithm with R code is provided to implement this method. The proposed method allows for high flexibility and clear targeting of desired effects, providing one solution to the bias-variance-generalizability tradeoff.
Keywords: Bias-variance tradeoff, causal inference, cluster-randomized trials, non-parametric estimation, permutation inference, quasi-experiments
@@ 1 Introduction
Staggered treatment adoption occurs in a wide variety of settings, including both observational and randomized contexts. In observational and quasi-experimental studies, panel data methods are commonly used to analyze the effect of a policy implementation or an exogenous shock. Stepped-wedge cluster-randomized trials have also become a common approach for the analysis of health, education, or other social policies, especially with a phased or gradual implementation. Across settings, the analysis of the data generated from staggered treatment settings requires careful consideration of the desired estimand, assumptions about the treatment effect, consideration of heterogeneity across units, time periods, and treatment regimens, and appropriate consideration of variance and correlation.
@@
This complexity has led to the development of a wide array of methods, commonly found in the econometrics literature surrounding panel data, difference-in-differences (DID), and staggered treatment adoption and in the biostatistics literature surrounding stepped-wedge trials (SWTs). Key developments in SWTs include approaches to interpret the targeted estimand (see, e.g., Twisk et al. (2016)), design-based considerations (see, e.g., Matthews & Forbes (2017) and Li et al. (2023)), robust inference (see, e.g., Wang & De Gruttola (2017), Hughes et al. (2020), Maleyeff et al. (2023)), and a variety of analytic approaches, discussed below. While development in both areas is still very much ongoing, recent reviews of the staggered adoption literature can be found in, among others, Baker et al. (2022), de Chaisemartin & D’Haultfœuille (2023), Roth et al. (2023), and Borusyak et al. (2024).
@@
What these new developments generally share is an acknowledgment of the need for careful selection of the target estimand and modeling of the treatment effect in order to estimate that estimand without bias. In the stepped-wedge setting, this is often done using a parametric or semi-parametric model for the treatment effect that can account for heterogeneity as in Hooper et al. (2016), Nickless et al. (2018), Kenny et al. (2022), Maleyeff et al. (2023), and Lee & Cheung (2024). It can, however, also be done using appropriate weighting of non-parametric estimators like those proposed by Thompson et al. (2018) and Kennedy‐Shaffer et al. (2020). In the staggered treatment adoption setting, approaches to address the bias that can arise in two-way fixed-effects models (see, e.g., Goodman-Bacon (2021) and Imai & Kim (2021)) have included adjusting the interpretation based on the identifying assumptions as in Athey & Imbens (2022), specifying “dynamic” (time-varying) treatment effects as in Sun & Abraham (2021), and restricting or combining with specified weights effect estimators targeting specific group-time effects as in Callaway & Sant’Anna (2021) and de Chaisemartin & D’Haultfœuille (2020). Lindner & McConnell (2021) highlight the relationship between these approaches, noting a similarity between the approaches of weighting specific effects in the DID and SWT settings.
@@
This paper introduces a generalized estimator that can target a variety of estimands specified by investigators under different sets of assumptions about treatment effect homogeneity. This class of estimators is constructed by taking weighted averages of simple two-by-two DID estimators, and finding the minimum-variance such weighting that targets the desired estimand. This allows one approach to be used across different assumption settings and different target estimands; it further allows different a priori variance assumptions to be considered and, if the assumption is correct, minimizes the corresponding variance of the estimator. This approach also eases interpretability by identifying the weights on the estimators as well as the weights on the individual observations and permits sensitivity analysis by assessing the estimand and relative efficiency under different settings.
@@
Section 2 of this paper describes the approach to constructing the estimator and its properties, with the full algorithm in Section 2.8. Section 3 provides a toy example to illuminate the algorithm in a few simple settings. Section 4 re-analyzes data from a SWT by Trajman et al. (2015) using this method. Section 5 comments on the advantages and limitations of the method, as well as future areas for research.
@@ 2.1 Notation
Consider a setting with $J$ periods ($j=1,\ldots,J$) and $N$ units of analysis ($i=1,\ldots,N$), which may be clusters or individual study units, for a total of $NJ$ observations. Denote by $Y_{ij}$ the outcome (or average outcome) in unit $i$ in period $j$ and by $X_{ij}$ the indicator of whether unit $i$ was treated/exposed in period $j$ (these are used interchangeably without, with a preference for “treated” for simplicity). Let $T_{i}=\min\{j:~{}X_{ij}=1\}$ be the time period in which unit $i$ was first treated. The staggered adoption assumption is made that once treated, a unit remains treated for the duration of the study. Note that de Chaisemartin & D’Haultfœuille (2020) consider more general settings with identifying assumptions that do not require staggered adoption; generalizations of the following methods to those settings may be feasible but are not considered here. Note also that there may be multiple units in the same sequence, referring to a pattern of treatment (under the staggered adoption assumption, this is fully specified by $T_{i}$).
Without loss of generality, the unit indices are ordered from the earliest treatment adopter to the latest. That is, if $i<i^{\prime}$, then $T_{i}\leq T_{i^{\prime}}$. Units with the same adoption time may be ordered in any way, as long as the ordering remains consistent throughout the notation.
@@ 2.2 Two-by-two DID estimators
These two-by-two DID estimators can be partitioned into six mutually exclusive categories based on the treatment pattern of both units. Under the staggered treatment assumption, since $j<j^{\prime}$, each cluster can either be untreated in both periods, treated in both periods, or untreated in period $j$ and treated in period $j^{\prime}$. Moreover, since $i<i^{\prime}$, $T_{i}\leq T_{i^{\prime}}$ so if unit $i$ is untreated in any period, unit $i^{\prime}$ must be as well. This leaves six possible combinations of treatment patterns, summarized in Table 1, along with the number in each category.
@@ 2.3 Expected value of two-by-two DID estimators
This paper focuses on additive treatment effects, and define the effect of treatment in period $j$ on unit $i$, which first adopted treatment in period $T_{i}$, by $\theta_{i,T_{i},j}$:
@@
where $Y_{ij}(1)$ is the potential outcome for unit $i$ in period $j$ if the unit is first treated in period $T_{i}\leq j$ and $Y_{ij}(0)$ is the potential outcome for unit $i$ in period $j$ if the unit is first treated after period $j$ or never treated. Note that multiplicative treatment effects can be considered by using a log transformation on the outcome. If the assumptions hold on the multiplicative scale, the estimator can then be computed using the log-transformed outcomes as the $Y_{ij}$ values. This has been discussed in staggered adoption settings (see, e.g., Kahn-Lang & Lang (2020) and Goodman-Bacon & Marcus (2020)) and in SWT settings (see, e.g., Kennedy‐Shaffer et al. (2020) and Kennedy-Shaffer & Lipsitch (2020)). To ensure consistency, the following no-spillover and no-anticipation assumptions are also needed; versions of these are common in both the staggered adoption and SWT literatures.
@@
Further simplification of this expectation depends on assumptions about the heterogeneity of the treatment effects by unit, time-on-treatment (i.e., exposure time), and period (i.e., calendar time). Many assumption settings are feasible; this paper considers the following five and defines a simplified treatment effect notation for each.
Assumption S 1
No additional assumptions: the treatment effect may vary by unit, exposure time, and/or calendar time. The treatment effect is denoted by $\theta^{(1)}_{i,j,a}$ in period $j$ for unit $i$ in its $a$th treated period. That is, $\theta_{i,T_{i},j}=\theta^{(1)}_{i,j,j-T_{i}+1}$ for all $i,j$ where $X_{ij}=1$.
Assumption S 2
Unit homogeneity: the treatment effect may vary by exposure time and/or calendar time, but does not vary by units in the same treatment adoption sequence. The treatment effect is denoted by $\theta^{(2)}_{j,a}$ in period $j$ for a unit in its $a$th treated period. That is, $\theta_{i,T_{i},j}=\theta^{(2)}_{j,j-T_{i}+1}$ for all $i,j$ where $X_{ij}=1$.
Assumption S 3
Unit and calendar-time homogeneity: the treatment effect may vary only by exposure time. The treatment effect is denoted by $\theta^{(3)}_{a}$ for a unit in its $a$th treated period. That is, $\theta_{i,T_{i},j}=\theta^{(3)}_{j-T_{i}+1}$ for all $i,j$ where $X_{ij}=1$.
Assumption S 4
Unit and exposure-time homogeneity: the treatment effect may vary only by calendar time. The treatment effect is denoted by $\theta^{(4)}_{j}$ in period $j$. That is, $\theta_{i,T_{i},j}=\theta^{(4)}_{j}$ for all $i,j$ where $X_{ij}=1$.
Assumption S 5
Full homogeneity: the treatment effect does not vary by unit, exposure time, or calendar time. The treatment effect is denoted by $\theta^{(5)}$. That is, $\theta_{i,T_{i},j}=\theta^{(5)}$ for all $i,j$ where $X_{ij}=1$.
@@ 2.4 Overall estimator form
Further, define $\bm{y}=\begin{pmatrix}Y_{1,1}&Y_{1,2}&\cdots&Y_{1,J}&Y_{2,1}&\cdots&Y_{N,J}\end{pmatrix}^{T}$ as the $NJ$-vector of observed outcomes ordered by unit $i$ and period $j$, respectively. Finally, let $\bm{A}$ be the $\binom{N}{2}\binom{J}{2}\times NJ$ matrix such that $\bm{d}=\bm{A}\bm{y}$. Note that each row of $\bm{A}$ corresponds to a unique two-by-two DID estimator and includes two entries of 1, two entries of -1, with remaining entries all 0. An algorithm to generate $\bm{A}$ for any $N$ and $J$ is given in Appendix A.
@@ 2.5 Unbiased estimation of target estimand
For any assumption setting, denote the vector of all unique treatment effects by $\bm{\theta}$, with a superscript to clarify the assumption setting if desired. For concreteness, order the treatment effects by cluster $i$, period $j$, and exposure time $a$, respectively, when necessary. Let $\theta_{e}$ be the desired estimand, which must be a linear combination of the unique treatment effects in $\bm{\theta}$. Then $\theta_{e}=\bm{v}^{T}\bm{\theta}$ for some vector $\bm{v}$ of weights on the unique treatment effects. For example, a vector $\bm{v}$ with all zero entries except for a 1 in one entry would pick out a single unique treatment effect, and a vector $\bm{v}$ with equal entries summing to 1 would average over all unique treatment effects. In addition to averages of certain effects, $\theta_{e}$ could also be a difference between two treatment effects, or a more complicated linear combination, as desired.
@@
For any assumption setting and vector of unique treatment effects $\bm{\theta}$ and any target estimand $\theta_{e}=\bm{v}^{T}\bm{\theta}$, $\hat{\theta}$ is an unbiased estimator of $\theta$ if $\bm{F}^{T}\bm{w}=\bm{v}$.
@@
$E[\hat{\theta}]$ $=E\left[\bm{w}^{T}\bm{d}\right]=\bm{w}^{T}E[\bm{d}]=\bm{w}^{T}F\bm{\theta}=\bm{v}^{T}\bm{\theta}=\theta_{e},$
as desired.
@@
Proof. The proof relies on the linear algebra results on non-homogeneous systems of linear equations and the rank-nullity theorem (George & Ajayakumar (2024), pp. 41 and 89, respectively), as well as the following two lemmata. The full proofs of the theorem and lemmata, are given in Appendix B.
Lemma 4
Let $\bm{F}$ and $\bm{A}$ be as defined previously. Then $ker(\bm{A}^{T})\subset ker(\bm{F}^{T})$.
Lemma 5
Let $\bm{A}$ be as defined previously for a setting with $N\geq 2$ units and $J\geq 2$ periods. Then $rank(A)=rank(A^{T})=(N-1)(J-1)$.

### v3

# A Generalized Difference-in-Differences Estimator for Randomized Stepped-Wedge and Observational Staggered Adoption Settings

@@ Abstract
Staggered treatment adoption arises in the evaluation of policy impact and implementation in many settings, including both randomized stepped-wedge trials and non-randomized quasi-experiments with panel data. In both settings, getting an interpretable, unbiased effect estimate requires careful consideration of the target estimand and possible treatment effect heterogeneities. This paper proposes a novel non-parametric approach to this estimation for either setting. By constructing an estimator using weighted averages of two-by-two difference-in-differences comparisons as building blocks, the investigator can target the desired estimand for any assumed treatment effect heterogeneities. This provides desirable bias and interpretation properties while using the comparisons efficiently to mitigate the loss of precision, without requiring correct variance specification. The methods are demonstrated for both a randomized stepped-wedge trial on the impact of novel tuberculosis diagnostic tools and an observational staggered adoption study on the effects of COVID-19 vaccine financial incentive lotteries in U.S. states; these are compared to analyses using previous methods. A full algorithm with R code is provided to implement this method and to compare against existing methods. The proposed method allows for high flexibility and clear targeting of desired effects, providing one solution to the bias-variance-generalizability tradeoff.
Keywords: Causal inference, cluster-randomized trials, natural experiments, panel data, quasi-experimental design
@@ 1 Introduction
Staggered treatment adoption occurs in a wide variety of settings, including both observational and randomized contexts. In observational and quasi-experimental studies, panel data methods are commonly used to analyze the effect of a policy implementation or an exogenous shock. Stepped-wedge cluster-randomized trials have also become a common approach for the analysis of health, education, or other social policies, especially those requiring a phased or gradual implementation. Across settings, the analysis of the data generated from staggered treatment adoption requires careful consideration of the desired estimand, assumptions about the treatment effect, consideration of heterogeneity across units, time periods, and treatment regimens, and appropriate consideration of variance and correlation.
@@
This complexity has led to the development of a wide array of methods, commonly found in somewhat distinct areas: (1) the econometrics literature surrounding panel data, difference-in-differences (DID), and staggered treatment adoption; and (2) the biostatistics literature surrounding stepped-wedge trials (SWTs). Key developments in SWTs include approaches to interpret the targeted estimand (see, e.g., Twisk et al. (2016)), design-based considerations (see, e.g., Matthews & Forbes (2017) and Li et al. (2024)), robust inference (see, e.g., Wang & De Gruttola (2017), Hughes et al. (2020), Maleyeff et al. (2023), and Wang, Wang & Li (2024)), and a variety of analytic approaches, discussed below. While development in both areas is still very much ongoing, recent reviews of the observational staggered adoption methods literature include, among others, Baker et al. (2022), de Chaisemartin & D’Haultfœuille (2023), Roth et al. (2023), and Borusyak et al. (2024). Review papers in specific disciplines also highlight how to interpret the assumptions and estimands therein; e.g., in health policy, see Feng & Bilinski (2024), Feng et al. (2024), and Wang, Hamad & White (2024).
These new developments highlight the need for careful selection of the target estimand and, for the most part, propose an identification strategy for unbiased estimation via modeling of the treatment effect in a regression framework. In the stepped-wedge setting, this is often done using a parametric or semi-parametric model for the treatment effect that can account for heterogeneity as in Hooper et al. (2016), Nickless et al. (2018), Kenny et al. (2022), Maleyeff et al. (2023), Lee & Cheung (2024), and Wang, Wang & Li (2024). It can, however, also be done using appropriate weighting of non-parametric estimators like those proposed by Thompson et al. (2018) and Kennedy‐Shaffer et al. (2020). In addition, the stepped-wedge literature has discussed the importance of using information efficiently, by combining “horizontal” (within-unit) and “vertical” (within-time) comparisons (see, e.g., Matthews & Forbes (2017) and Kasza, Taljaard & Forbes (2019)).
@@
In the observational staggered adoption setting, approaches to address the biases that can arise in two-way fixed-effects models (see, e.g., Goodman-Bacon (2021) and Imai & Kim (2021)) have included adjusting the interpretation based on the identifying assumptions as in Athey & Imbens (2022), specifying “dynamic” (time-varying) treatment effects as in Sun & Abraham (2021), and isolating, excluding, or weighting effect estimators targeting specific group-time effects as in Callaway & Sant’Anna (2021) and de Chaisemartin & D’Haultfœuille (2020). Lindner & McConnell (2021) highlight the relationship between these approaches, noting a similarity between the approaches of weighting specific effects in the DID and SWT settings.
@@
This paper introduces a generalized estimator framework that can target a variety of estimands specified by investigators under different sets of assumptions about treatment effect homogeneity. This class of estimators is constructed by taking weighted averages of simple two-by-two DID estimators, and finding the minimum-variance weighting that targets the desired estimand. The framework has the advantages of: (1) unifying methods across different settings, assumptions, and target estimands; (2) minimizing variance under certain variance-specific assumptions while preserving validity; (3) easing interpretability and improving robustness by explicitly identifying weights on the two-by-two DID estimators and on the observations themselves; and (4) permitting sensitivity analysis across heterogeneity and variance assumptions. Several existing staggered adoption estimators are special cases of this approach, allowing easy comparisons across methods.
@@
Section 2 of this paper describes the approach to constructing the estimator and its properties, with the full algorithm in Section 2.8. Section 3 provides a toy example to illuminate the algorithm in a few simple settings. Section 4 situates several existing estimators as special cases of this framework and the assumptions they use, while noting others which do not fit within the framework. Section 5 re-analyzes data from a SWT on the effect of improved diagnostic methods on tuberculosis outcomes by Trajman et al. (2015) using this method. Section 6 re-analyzes data from an observational staggered adoption setting described by Fuller et al. (2022) on the effects of financial incentive lotteries to encourage uptake of COVID-19 vaccines in the United States. For both examples, comparisons are made to the results using existing methods. Section 8 comments on the advantages and limitations of the method, as well as future areas for research.
@@ 2.1 Notation
Consider a setting with $J$ periods ($j=1,\ldots,J$) and $N$ units of analysis ($i=1,\ldots,N$), which may be clusters or individual study units, for a total of $NJ$ observations. Denote by $Y_{ij}$ the outcome (or average outcome) in unit $i$ in period $j$ and by $X_{ij}$ the indicator of whether unit $i$ was treated/exposed in period $j$ (these are used interchangeably throughout, with a preference for “treated” for simplicity). Let $T_{i}=\min\{j:~{}X_{ij}=1\}$ be the time period in which unit $i$ was first treated. The staggered adoption assumption is made that once treated, a unit remains treated for the duration of the study. Note that de Chaisemartin & D’Haultfœuille (2020) consider more general settings with identifying assumptions that do not require staggered adoption; generalizations of the following methods to those settings may be feasible but are not considered here. Note also that there may be multiple units in the same sequence, referring to a pattern of treatment (under the staggered adoption assumption, this is fully specified by $T_{i}$).
Without loss of generality, the unit indices are ordered from the earliest treatment adopter to the latest. That is, if $i<i^{\prime}$, then $T_{i}\leq T_{i^{\prime}}$. Units with the same adoption time may be ordered in any way, as long as the ordering remains consistent.
@@ 2.2 Two-by-two DID estimators
These two-by-two DID estimators can be partitioned into six mutually exclusive categories based on the treatment pattern of both units. Under the staggered treatment assumption, since $j<j^{\prime}$, each cluster can either be untreated in both periods, treated in both periods, or untreated in period $j$ and treated in period $j^{\prime}$. Moreover, since $i<i^{\prime}$, $T_{i}\leq T_{i^{\prime}}$, so there are no periods where unit $i$ is untreated and unit $i^{\prime}$ is treated. This leaves six possible combinations of treatment patterns, summarized in Table 1, along with the number in each category.
@@ 2.3 Expected value of two-by-two DID estimators
This paper focuses on additive treatment effects, so the effect of treatment in period $j$ on unit $i$, which first adopted treatment in period $T_{i}$, denoted $\theta_{i,T_{i},j}$, is given by:
@@
where $Y_{ij}(1)$ is the potential outcome for unit $i$ in period $j$ if the unit is first treated in period $T_{i}\leq j$ and $Y_{ij}(0)$ is the potential outcome for unit $i$ in period $j$ if the unit is first treated after period $j$ or never treated. Note that multiplicative treatment effects can be considered by using a log transformation on the outcome. If the assumptions hold on the multiplicative scale, the estimator can then be computed using the log-transformed outcomes as the $Y_{ij}$ values. This has been discussed in staggered adoption settings (see, e.g., Kahn-Lang & Lang (2020) and Goodman-Bacon & Marcus (2020)) and in SWT settings (see, e.g., Kennedy‐Shaffer et al. (2020) and Kennedy-Shaffer & Lipsitch (2020)).
To ensure unbiased estimation, the following assumptions (no spillover and no anticipation, respectively) are also needed; versions of these are common in both the staggered adoption and SWT literatures.
@@
Further simplification of this expectation depends on assumptions about the heterogeneity of the treatment effects by unit, time-on-treatment (i.e., exposure time), and period (i.e., calendar time). Many assumption settings are feasible; this paper considers five and defines a simplified treatment effect notation for each; these are described in Table 2.
Table 2: Treatment effect heterogeneity assumption settings.
Assumed
homogeneities
Allowed heterogeneities Notation Effect values
S1 None Unit ($i$)
Exposure time ($a$)
Calendar time ($j$)
$\theta^{(1)}_{i,j,a}$ $\theta_{i,T_{i},j}=\theta^{(1)}_{i,j,j-T_{i}+1}$
S2 Unit Exposure time ($a$)
Calendar time ($j$)
$\theta^{(2)}_{j,a}$ $\theta_{i,T_{i},j}=\theta^{(2)}_{j,j-T_{i}+1}$
S3 Unit, calendar time Exposure time ($a$) $\theta^{(3)}_{a}$ $\theta_{i,T_{i},j}=\theta^{(3)}_{j-T_{i}+1}$
S4 Unit, exposure time Calendar time ($j$) $\theta^{(4)}_{j}$ $\theta_{i,T_{i},j}=\theta^{(4)}_{j}$
S5 Unit, exposure time,
calendar time
None $\theta^{(5)}$ $\theta_{i,T_{i},j}=\theta^{(5)}$
@@ 2.4 Overall estimator form
Further, define $\bm{y}=\begin{pmatrix}Y_{1,1}&Y_{1,2}&\cdots&Y_{1,J}&Y_{2,1}&\cdots&Y_{N,J}\end{pmatrix}^{T}$ as the $NJ$-vector of observed outcomes ordered by unit $i$ and period $j$, respectively. Finally, let $\bm{A}$ be the $\binom{N}{2}\binom{J}{2}\times NJ$ matrix such that $\bm{d}=\bm{A}\bm{y}$. Note that each row of $\bm{A}$ corresponds to a unique two-by-two DID estimator and includes two entries of 1 and two entries of -1, with the remaining entries all 0. An algorithm to generate $\bm{A}$ for any $N$ and $J$ is given in Appendix A.
@@ 2.5 Unbiased estimation of target estimand
For any assumption setting, denote the vector of all unique treatment effects by $\bm{\theta}$, with a superscript to clarify the assumption setting if desired. For concreteness, order the treatment effects by cluster $i$, period $j$, and exposure time $a$, respectively, when necessary. Let $\theta_{e}$ be the desired estimand, which must be a linear combination of the unique treatment effects in $\bm{\theta}$. Then $\theta_{e}=\bm{v}^{T}\bm{\theta}$ for some vector $\bm{v}$ of weights on the unique treatment effects. For example, a vector $\bm{v}$ with all zero entries except for a 1 in one entry would pick out a single unique treatment effect, and a vector $\bm{v}$ with equal entries summing to 1 would average over all unique treatment effects. In addition to averages of certain effects, $\theta_{e}$ could also be a difference between two treatment effects, or a more complicated linear combination, as desired. For example, heterogeneity of effects could be assessed by estimating the difference between two effects as discussed by Shahn & Hatfield (2024) and Xu et al. (2024).
@@
For any assumption setting with corresponding vector of unique treatment effects $\bm{\theta}$ and any target estimand $\theta_{e}=\bm{v}^{T}\bm{\theta}$, assuming that $\bm{w}$ is independent of $\bm{y}$, $\hat{\theta}$ is an unbiased estimator of $\theta_{e}$ if $\bm{F}^{T}\bm{w}=\bm{v}$.
@@
$E[\hat{\theta}]$ $=E\left[\bm{w}^{T}\bm{d}\right]=\bm{w}^{T}E[\bm{d}]=\bm{w}^{T}F\bm{\theta}=\bm{v}^{T}\bm{\theta}=\theta_{e}.$ (6)
@@
Proof. The proof relies on linear algebra results on non-homogeneous systems of linear equations and the rank-nullity theorem (George & Ajayakumar (2024), pp. 41 and 89, respectively). The full proof of the theorem, with two necessary lemmata, is given in Appendix B.